A straight, horizontal wire of mass 10 mg and length 1.0 m carries a current of 2.0 A. What minimum magnetic field B should be applied in the region so that the magnetic force on the wire may balance its weight?

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. Mass = 10 mg = 10^–5 kg Length = 1 m I = 2 A, B = ? Now, Mg = ilB ⇒B = mg/il = 10^-5 *9.8/2 * 1 = 4.9 * 10^-5 T